OpenStudy (anonymous):

6.02] For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + y = 8 −x − y = −5 8 + 2x 8 − 2x −8 − 2x −8 + 2x

OpenStudy (anonymous):

to isolate y is to put it on one side of the equation by itself. try that.

OpenStudy (anonymous):

Uh, I don't know, c?

OpenStudy (anonymous):

2x + y = 8 take this equation and rearrange it so y is on the side by itself

OpenStudy (anonymous):

I'm confused. How do u do that?

OpenStudy (anonymous):

2x+y=8 2x+y-2x=8-2x y=8-2x

OpenStudy (anonymous):

take 2x away from both sides. The equation is still legit and it works :)

OpenStudy (anonymous):

So d?

OpenStudy (anonymous):

b :)

OpenStudy (anonymous):

Ok, thx so much! :D I don't understand this!! happy thanksgiving!!